### v2

# LangBridge: Interpreting Image as a Combination of Language Embeddings

Abstract
Recent years have witnessed remarkable advances in Large Vision-Language Models (LVLMs), which have achieved human-level performance across various complex vision-language tasks. Following LLaVA’s paradigm, mainstream LVLMs typically employ a shallow MLP for visual-language alignment through a two-stage training process: pretraining for cross-modal alignment followed by instruction tuning. While this approach has proven effective, the underlying mechanisms of how MLPs bridge the modality gap remain poorly understood. Although some research has explored how LLMs process transformed visual tokens, few studies have investigated the fundamental alignment mechanism. Furthermore, the MLP adapter requires retraining whenever switching LLM backbones. To address these limitations, we first investigate the working principles of MLP adapters and discover that they learn to project visual embeddings into subspaces spanned by corresponding text embeddings progressively. Based on this insight, we propose LangBridge, a novel adapter that explicitly maps visual tokens to linear combinations of LLM vocabulary embeddings. This innovative design enables pretraining-free adapter transfer across different LLMs while maintaining performance. Our experimental results demonstrate that a LangBridge adapter pre-trained on Qwen2-0.5B can be directly applied to larger models such as LLaMA3-8B or Qwen2.5-14B while maintaining competitive performance. Overall, LangBridge enables interpretable vision-language alignment by grounding visual representations in LLM vocab embedding, while its plug-and-play design ensures efficient reuse across multiple LLMs with nearly no performance degradation. See our project page at https://jiaqiliao77.github.io/LangBridge.github.io/.
Equal contribution.
1 Introduction
Figure 1: Comparison of different connector types in LVLM: (a) The MLP directly maps visual features into the LLM’s text embedding space. (b) The Ovis method uses a visual embedding table to produce structural visual embeddings and align the modalities. (c) LangBridge decomposes visual features into weighted combinations of LLM’s vocabulary vectors to form the visual embeddings
Recent years have witnessed remarkable progress in Large Vision-Language Models (LVLMs) [1, 18, 53, 21, 15, 5, 39, 17], which have demonstrated impressive capabilities in complex vision-language tasks, showcasing sophisticated understanding and reasoning across both visual and textual modalities. As a result, these models are increasingly being deployed in real-world applications, ranging from autonomous driving to robotics [31, 3, 24].
As shown in Figure 1 (a), mainstream Large Vision-Language Models follow the architecture of LLaVA [18], leveraging Vision Transformers (ViTs) for visual feature extraction and Multi-Layer Perceptrons (MLPs) to transform these features into visual embeddings aligned with the LLM’s text embedding dimension. These models also adopt LLaVA’s two-stage training process: (1) MLP warm-up for cross-modal alignment and (2) full-model instruction tuning. While this approach has proven effective, two critical challenges remain: 1) the underlying mechanisms of how MLPs bridge the modality gap are poorly understood. Although some research has explored how LLMs process transformed visual tokens [12, 49], few studies have investigated the fundamental alignment mechanisms of MLPs, and 2) the MLP adapter requires retraining whenever switching LLM backbones due to input dimension mismatch and distributional shifts in feature representations.
To address these limitations, we systematically investigate the working principles of MLP adapters through two critical aspects: 1) how visual semantics are encoded in transformed embeddings, and 2) how MLPs learn cross-modal alignment during training. Our methodology employs visual analysis across four distinct training stages: 1) Pretrain-100 steps, 2) Pretrain-1000 steps, 3) Pretrain-2000 steps, and 4) SFT-tuned Model. For each stage, we extract visual features using Vision Transformers, transform them through stage-specific MLPs into visual embeddings, and compute cosine similarities between these visual embeddings and text embeddings. Through this process, we identify the top semantically related text tokens for each visual embedding at each training phase.
As visualized in Figure 2 through circular graphs where nodes represent different training phases spanning from early to late stages (e.g. 1- 4), we can draw two key conclusions: 1. Visual embeddings demonstrate strong correlations with semantically-related text tokens. As shown in Figure 2, green apple image patch exhibits high similarity with text tokens such as “Green Apple”. Similarly, the silhouetted sunset scene correlates with concepts like “Eclipse Sun”. This suggests that MLPs effectively map visual features into the text embedding space of LLMs, particularly in regions close to the semantic spaces formed by the corresponding text embeddings. 2. The projection capability of MLPs develops progressively during training. Figure 2 illustrates this evolution through multiple semantic pathways: from “numerous” to ”five”, from ”Arts” to “Green Apple”, and in the right scene, from “We Plat” to “Shakespeare”, from “Kiss” to “Wed Kiss”. These relationships become increasingly well-defined through the training stages, confirming our hypothesis that MLPs gradually learn to map visual features into regions proximate to their corresponding text embedding subspaces.
Building upon this insight, we introduce an explicit transformation approach called Language Basis Vector Projection. This approach explicitly projects visual features into the text embedding subspace of the LLM by representing them as linear combinations of the LLM’s vocabulary embeddings. Based on this theoretical foundation, we propose LangBridge, a novel adapter architecture. The adapter first maps visual features to probability distributions that capture their semantic similarity with vocabulary text embeddings in the LLM. Subsequently, it generates the final visual embeddings by combining these text embeddings, weighted according to the computed probabilities. This innovative design elegantly addresses the challenge of model retraining when switching between different LLM backbones. The key insight is that our adapter only learns the linear combination relationship between visual patches and vocabulary embeddings, rather than directly mapping between different embedding dimensions. Since the adapter’s output is always a probability distribution over a fixed vocabulary, it can be used to weigh and combine the corresponding vocabulary embeddings of any target LLM, regardless of their embedding dimensions.
Extensive experiments across multiple benchmarks and LLMs demonstrate the effectiveness of our approach. A key advantage of LangBridge is its ability to enable pre-training-free transfer between different LLMs. Specifically, LangBridge trained on smaller models (e.g., Qwen2-0.5B or LLaMA-3-8B) can be directly applied to larger models like Qwen2-7B and Qwen2.5-7B during supervised fine-tuning while maintaining comparable performance. This capability significantly reduces computational costs by eliminating the need for repeated pre-training when switching between different LLMs. Our key contributions can be summarized as follows:
  1. 1.
    We investigate the working principles of MLP adapters, with a particular focus on the semantic meaning of visual embeddings and how MLPs learn cross-modal alignment during training.
  2. 2.
    LangBridge: A Novel Adapter for LVLMs. We introduce LangBridge, a novel adapter that transforms the visual features into visual embeddings by decomposing them into linear combinations of the LLM’s vocabulary text embeddings.
  3. 3.
    LangBridge enables pre-training-free reuse between different large language models, significantly reducing the cost of pretraining.
Figure 2: Progressive semantic alignment in MLP adapters across training stages. Circular graphs demonstrate the evolution of visual-text token associations through four training phases (Pretrain-100, Pretrain-1000, Pretrain-2000, and Final-SFT). (a) For a bunch of apples image, the MLP progressively refines associations from the meaningless text (“numerous”, “Arts”) to meaningful text (“five”, “Green Apple”). (b) For a sunset silhouette scene and poem, semantic mappings evolve from the meaningless text (“We Plat”, “Kiss”) to contextually relevant tokens (“Shakespeare”, “Wed Kiss”), illustrating the MLP’s increasing capability to project visual features into LLM’s text embedding space.
2 Related Work
2.1 Visual encoders in LVLMs
In recent years, large language models (LLMs) [50, 23, 25] have achieved remarkable advances in natural language processing. However, their inability to directly process visual information limits their effectiveness in multimodal tasks. To overcome this limitation, visual encoders [28, 47, 34, 27, 29] have been integrated, serving as the “eyes” of large vision-language models to leverage LLMs’ extensive knowledge [45, 26, 38] and strong instruction-following capabilities. Visual models that learn representations using only visual signals, such as DINOv2 [27] and diffusion models [29], often face challenges in achieving high performance, while models like CLIP [28] and SigLIP [47] align visual and text modalities effectively by training on extensive image-text datasets using contrastive learning. Recent research [41, 42, 36, 35] has also advocated for combining multiple visual encoders to better capture complex visual information. Nonetheless, aligning visual encoders with large language models remains challenging [44] due to differences in their representational spaces.
2.2 Connectors in LVLMs
To merge the visual extraction strengths of visual encoders with LLMs’ broad knowledge and instruction-following skills, researchers have explored methods to connect these components [54, 37, 43, 33, 48, 8, 9, 22, 4, 51, 6, 2, 10, 16, 52]. Pioneering work represented by models such as Flamingo [1], which uses cross-attention mechanisms, and BLIP-2 [13], which introduces the Q-former architecture, marks significant progress in this field. Subsequently, LLaVA [19, 18] adopted a two-stage learning approach, using a multilayer perceptron to map visual information into language embedding space, establishing itself as a most widely applicable alignment method [53, 21, 15, 5]. Departing from previous approaches, we propose an innovative mapping technique. Rather than mapping visual information directly to language embedding space, we introduce “language basis vectors” to learn linear combination weights of these basis vectors, creating a new way to represent visual information.
3 How LVLMs process the visual information
In this section, we first introduce the preliminaries of current LVLMs (Sec 3.1). Then we conduct a comprehensive analysis of how MLPs bridge the modality gap through two aspects: 1. investigating how visual semantics are encoded in transformed embeddings (Sec 3.2), and 2. exploring how MLPs learn cross-modal alignment during training (Sec 3.3). Based on these insights, we propose the Language Basis Vector Projection Method (Sec 3.4).
3.1 Preliminaries
As illustrated in Figure 1 (a), mainstream LVLMs follow LLaVA’s [18] architecture with three key components: (1) Visual Encoder: Vision Transformers (ViTs) that extract visual features from input images; (2) Connector: Multilayer Perceptrons (MLPs) that transform these visual features into the LLM’s embedding space; and (3) Large Language Model (LLM): The language model that processes the aligned embeddings and generates responses.
3.2 Semantic Analysis of Visual Embeddings
To understand how visual semantics are encoded in transformed embeddings, we analyze the relationship between visual embeddings and text embeddings through cosine similarity:
$\text{Similarity}(v_{i},t_{k})=\frac{\mathbf{v}_{i}\cdot\mathbf{t}_{k}}{\|\mathbf{v}_{i}\|\|\mathbf{t}_{k}\|}$ (1)
where $\mathbf{v}_{i}$ represents the visual embedding and $\mathbf{t}_{k}$ denotes the LLM’s vocabulary text embedding. As shown in Figure 2, visual embeddings exhibit strong correlations with semantically related text tokens. For example, image patches of red apples show high similarity to tokens like ‘red’ and ‘Apple’. Moreover, some image tokens convey global information, for instance, ‘5,’ which refers to the number of apples. This result demonstrates that visual embeddings are within the LLM’s text embedding space through the MLP mapping, and are particularly close to their semantically related text embeddings.
3.3 Cross-Modal Alignment Learning Process
To investigate the development of cross-modal alignment capabilities, we conduct a temporal analysis across different training stages: Pretrain-100 steps, Pretrain-1000 steps, Pretrain-2000 steps, and final SFT. Our analysis reveals that MLPs undergo a progressive learning process in which they: 1) Initially establish basic visual-textual correspondences, 2) Gradually refine the projection of visual features into relevant text embedding subspaces, 3) Finally achieve robust semantic alignment between modalities. This evolutionary process is clearly demonstrated in Figure 2, where we observe increasing semantic alignment strength and precision as training progresses. These results support our hypothesis that MLPs gradually develop their projection capabilities, incrementally learning to map visual features into regions close to the subspace spanned by their corresponding text embeddings.
3.4 Language Basis Vector Projection Method
Based on our findings that MLPs learn to project visual features into regions close to the subspace spanned by their corresponding text embeddings, we propose Language Basis Vector Projection as an explicit transformation approach. This method represents visual embeddings as weighted combinations of LLM vocabulary embeddings:
$\mathbf{v}=\sum_{k=1}^{N}\beta_{k}\mathbf{t}_{k}$ (2)
where $\beta_{k}$ represents the probability distribution over vocabulary tokens. This approach offers two key advantages: 1) it explicitly models the semantic relationship between visual and textual modalities, and 2) it enables pre-training-free transfer between different LLMs since it only learns the combination weights of Language Basis Vector rather than direct dimension mapping.
4 LangBridge
Figure 3: Overview of LangBridge architecture and workflow. (a) The architecture of LangBridge: LangBridge first extracts visual features through a Vision Encoder, then transforms them into visual embeddings by decomposing them into linear combinations of LLM’s vocabulary embeddings. These visual embeddings are then concatenated with text embeddings for LLM processing. (b) Linear Combination: LangBridge generates probability distributions over the vocabulary and multiplies them with text embeddings to form visual embeddings. (c) Vocabulary Embedding Selection: A subset of shared vocabulary embeddings is selected to reduce parameter count and optimization complexity while enabling cross-LLM reuse. (d) Adapter Reuse: LangBridge shows cross-LLM adaptability, allowing an adapter pre-trained on one LLM to be reused on a different LLM during the SFT stage.
In this section, we introduce LangBridge, a novel adapter that translates visual features into the language model’s embedding space based on the Language Basis Vector Projection Theory. Our key insight is to decompose visual features into linear combinations of the LLM’s vocabulary embeddings, enabling seamless integration with different LLM architectures. First, we present the core architecture of LangBridge (Sec 4.1), which operates by extracting visual features from images and projecting them into probability distributions over the LLM’s vocabulary space. These distributions are then used to construct visual tokens as weighted combinations of text embeddings. Second, to address the computational challenges posed by large vocabulary embedding tables (150,000 tokens), we introduce an efficient vocabulary selection strategy (Sec 4.2). By carefully selecting a subset of approximately 20,000 most frequent tokens shared across different LLMs, we significantly reduce the parameter size and enable reuse across different LLMs. Finally, we explain how the pretrained LangBridge can be seamlessly reused across different LLM architectures without retraining (Sec 4.3).
4.1 Architecture of LangBridge
We present the architecture of LangBridge, a novel vision-language adapter based on the Language Basis Vector Projection Theory. As illustrated in Figure 3, the framework operates through three stages:
Stage 1: Visual Feature Extraction. Given an input image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, we employ a Vision Transformer to extract patch-aligned visual features:
$\{v_{i}\}_{i=1}^{N}=\text{ViT}(\mathcal{I}),\quad v_{i}\in\mathbb{R}^{D}$ (3)
where $N=\frac{HW}{P^{2}}$ denotes the number of image patches with patch size $P$.
Stage 2: Probability Computation. After acquiring the visual features, we transform them into visual tokens and align them with the LLM’s text embedding space. This process is carried out by decomposing the visual features into linear combinations of the LLM’s vocabulary embeddings. As shown in Eq. 4, we proceed as follows: We first apply a two-layer MLP to project the visual features into the LLM’s text embedding space. Let $\mathbf{v}$ represent the visual feature, The projection is given by: $\mathbf{v}_{\text{proj}}=\text{MLP}(\mathbf{v})$, where $\mathbf{v}_{\text{proj}}\in\mathbb{R}^{D}$, and $D$ is the dimensionality of the text embeddings. Next, we append a linear layer $\mathbf{W}\in\mathbb{R}^{T\times D}$, where $T$ is the vocabulary size of the LLM, to produce the probability distribution $\mathbf{p}$, matching the LLM’s vocab size.
$\mathbf{p}=\mathbf{W}\cdot\text{MLP}(\mathbf{v})$ (4)
Stage 3: Linear combination of text embeddings. Once we have the probability distribution $\mathbf{p}$ from Stage 2, the visual features are represented as a weighted combination of the LLM’s text embeddings. The probability distribution $\mathbf{p}$ serves as the coefficients for a linear combination of the LLM’s vocabulary embeddings $\mathbf{e}_{i}$, where each $\mathbf{e}_{i}\in\mathbb{R}^{D}$ is a text embedding of the $i$-th token in the vocabulary.
Specifically, the visual tokens $\mathbf{v}_{\text{tokens}}$ are obtained as follows:
$\mathbf{v}_{\text{tokens}}=\sum_{i=1}^{T}p_{i}\mathbf{e}_{i}$
where $p_{i}$ is the $i$-th component of the probability distribution $\mathbf{p}$, and $T$ is the vocabulary size of the LLM. The resulting vector $\mathbf{v}_{\text{tokens}}\in\mathbb{R}^{D}$ represents the final visual token, which aligns with the LLM’s text embedding space.
4.2 Vocabulary Selection
A straightforward implementation of LangBridge would use the entire vocabulary embedding matrix of the LLM, which can be extremely large (nearly 1B size). To mitigate these issues, we select a compact subset of $N<20\text{k}$ tokens. To ensure the generalizability of the vocabulary and mitigate the parameter burden of LangBridge for efficient training, we merged the vocabularies of LLaMA and Qwen, retaining only the tokens present in both. Subsequently, we performed tokenization on the ShareGPT4V and LLaVA-CC3M-Pretrain-595K datasets to generate a vocabulary frequency table. From this table, we selected the top 19,200 most frequent tokens to construct our reduced vocabulary, thereby decreasing the vocabulary size. Ultimately, the word embeddings corresponding to these 19,200 vocabulary tokens are utilized as our language basis vectors.
4.3 Adapter Reuse Across Different LLMs
A key feature of LangBridge is its ability to be seamlessly reused across different LLM architectures after being pretrained. As shown in Figure 3(d), this reuse mechanism consists of two stages:
Stage1: Pretrain. During the pretraining stage, we first extract image features through the Vision Encoder, then use LangBridge to learn to map these features to the space of Shared Vocab Embedding of LLM-1, optimizing by the caption loss. In this stage, LangBridge outputs probability distributions over the selected vocabulary.
Stage2: SFT. When constructing a LVLM based on the new LLM-2, we do not need to retrain LangBridge in stage 1. Instead, we can directly reuse the adapters that were co-trained with LLM-1 in stage 1. The core principle behind this is that LangBridge only needs to out the probability distributions over the corresponding vocab instead of direct transform. When reusing LangBridge with different LLMs, such as Qwen2-0.5B and LLaMA3-8B, the key difficulty lies in the vocabulary (due to differences in vocabulary size). To enable reuse across different LLMs, as described in Figure 3 (c) and Section 4.2, we construct a shared vocabulary $V_{shared}$. The reusing process can be expressed as:
$P=\text{LangBridge}_{\text{LLM}_{1}}(I)\in\mathbb{R}^{|V_{\text{shared}}|}$ (5)
$\text{Visiontoken}_{\text{LLM}2}=P\cdot V_{\text{shared}}$ (6)
Here, $P$ represents the probability distribution over the shared vocabulary $V_{\text{shared}}$. The operation $P\cdot V_{\text{shared}}$ projects this distribution onto LLM-2’s vocabulary, resulting in the vision token $\text{Visiontoken}_{\text{LLM-2}}$, which is then sent to LLM-2 for processing.
5 Experiments
Table 1: Same-architecture transfer results. We evaluate the transferability of LangBridge between LLMs of the same architecture but different parameter sizes by using pretrained and fine-tuned connectors from Qwen2-0.5B (denoted as Qwen2-0.5B-Pretrain-LangBridge and Qwen2-0.5B-SFT-LangBridge, respectively). Red and blue percentages indicate performance improvement and degradation relative to baseline models, respectively.
SFT-LLM Connector Benchmark Results
GQA TextVQA MME MMBench MMVeT POPE SciQA
Qwen2-7B Qwen2-7B-Pretrain-MLPs 62.92 57.24 1938 72.7 35.5 87.8 79.44
- Qwen2-0.5B-Pretrain-LangBridge 63.03 (100.17%) 57.25 (100.02%) 1886 (97.32%) 71.7 (98.62%) 34.1 (96.06%) 88.2 (100.46%) 79.23 (99.74%)
- Qwen2-0.5B-SFT-LangBridge 63.15 (100.37%) 57.34 (100.17%) 1904 (98.25%) 71.0 (97.66%) 31.6 (89.01%) 88.3 (100.57%) 79.25 (99.76%)
Qwen2.5-7B Qwen2.5-7B-Pretrain-MLPs 62.70 57.83 1939 73.8 34.8 88.4 79.11
- Qwen2-0.5B-Pretrain-LangBridge 62.68 (99.97%) 57.35 (99.17%) 1871 (96.49%) 72.3 (97.97%) 34.7 (99.71%) 88.1 (99.66%) 81.23 (102.68%)
- Qwen2-0.5B-SFT-LangBridge 62.69 (99.98%) 57.94 (100.19%) 1915 (98.76%) 72.7 (98.51%) 35.6 (102.30%) 88.1 (99.66%) 77.1 (97.46%)
Qwen2.5-14B Qwen2.5-14B-Pretrain-MLPs 63.71 61.32 2038 78.2 37.7 88.1 85.59
- Qwen2-0.5B-Pretrain-LangBridge 63.75 (100.06%) 61.57 (100.41%) 1963 (96.32%) 76.2 (97.44%) 35.9 (95.23%) 88.2 (100.11%) 84.74 (99.01%)
- Qwen2-0.5B-SFT-LangBridge 63.92 (100.33%) 62.02 (101.14%) 1990 (97.64%) 77.4 (98.98%) 38.4 (101.86%) 87.6 (99.43%) 84.77 (99.04%)
Table 2: Cross-architecture transfer results. We evaluate the transferability of LangBridge between different LLM architectures by using pre-trained and fine-tuned connectors from LLaMA3-8B with Qwen2-7B, Qwen2.5-7B/14B, and Qwen2-0.5B with LLaMA3-8B.
SFT-LLM Connector Benchmark Results
GQA TextVQA MME MMBench MMVeT POPE SciQA
LLaMA3-8B LLaMA3-8B-Pretrain-MLPs 63.24 55.66 1736 71.0 31.0 87.1 78.61
- Qwen2-0.5B-Pretrain-LangBridge 64.14 (101.42%) 56.70 (101.87%) 1772 (102.07%) 70.1 (98.73%) 34.0 (109.68%) 87.5 (100.46%) 78.78 (100.22%)
- Qwen2-0.5B-SFT-LangBridge 63.96 (101.14%) 57.35 (103.04%) 1804 (103.92%) 69.4 (97.75%) 32.7 (105.48%) 87.5 (100.46%) 78.33 (99.64%)
Qwen2-7B Qwen2-7B-Pretrain-MLPs 62.92 57.24 1938 72.7 35.5 87.8 79.44
- LLaMA3-8B-Pretrain-LangBridge 62.90 (99.97%) 57.23 (99.98%) 1874 (96.70%) 72.1 (99.17%) 34.5 (97.18%) 87.6 (99.77%) 80.17 (100.92%)
- LLaMA3-8B-SFT-LangBridge 62.77 (99.76%) 57.08 (99.72%) 1915 (98.81%) 71.7 (98.62%) 33.2 (93.52%) 88.2 (100.46%) 78.83 (99.23%)
Qwen2.5-7B Qwen2.5-7B-Pretrain-MLPs 62.70 57.83 1939 73.8 34.8 88.4 79.11
- LLaMA3-8B-Pretrain-LangBridge 62.72 (100.03%) 57.78 (99.91%) 1878 (96.85%) 71.9 (97.43%) 32.3 (92.82%) 88.4 (100.00%) 80.59 (101.87%)
- LLaMA3-8B-SFT-LangBridge 63.03 (100.53%) 57.72 (99.81%) 1909 (98.45%) 71.1 (96.34%) 37.7 (108.33%) 90.1 (101.92%) 79.84 (100.92%)
Qwen2.5-14B Qwen2.5-14B-Pretrain-MLPs 63.71 61.32 2038 78.2 37.7 88.1 85.59
- LLaMA3-8B-Pretrain-LangBridge 63.88 (100.27%) 61.39 (99.89%) 2005 (98.38%) 77.1 (98.59%) 36.9 (97.88%) 87.9 (99.77%) 85.03 (99.35%)
- LLaMA3-8B-SFT-LangBridge 64.47 (101.19%) 61.35 (99.95%) 1980 (97.15%) 76.3 (97.57%) 38.3 (101.59%) 87.8 (99.66%) 84.23 (98.41%)
Table 3: Performance comparison of different paradigms. We compare our LangBridge with standard MLPs across seven widely used benchmarks. The results demonstrate that LangBridge achieves performance comparable to that of the standard MLP adapter.
LLM Connector Benchmark Results
GQA TextVQA MME MMBench MMVeT POPE SciQA
LLaMA3-8B MLPs 63.24 55.66 1736 71.0 31.0 87.1 78.61
- LangBridge 64.00 (101.20%) 56.55 (101.60%) 1775 (102.25%) 72.6 (102.25%) 31.1 (100.32%) 87.7 (100.69%) 77.98 (99.20%)
Qwen2.5-7B MLPs 62.70 57.83 1939 73.8 34.8 88.4 79.11
- LangBridge 62.88 (100.29%) 57.81 (99.97%) 1874 (96.65%) 73.8 (100.00%) 36.7 (105.46%) 88.3 (99.89%) 79.18 (100.09%)
In this section, we first introduce the experimental setup of this study in detail. Subsequently, we systematically compare the performance of the proposed LangBridge with baseline models across multiple evaluation metrics and highlight its unique advantage of enabling pre-training-free transfer between different large language models. Specifically, we explore two scenarios: same-architecture transfer and cross-architecture transfer, to demonstrate the unique advantage of the proposed paradigm fully. As a novel training paradigm, we also compare LangBridge’s performance with LLaVA on the same LLM (i.e., without transfer). Finally, we conduct an in-depth ablation study and provide a detailed analysis of the obtained results.
5.1 Experiment setup
5.1.1 Implementation Details.
Our architecture is composed of four key components: language model, visual encoder, connector, and language base vector. We selected LLaMA3-8B-Instruct [7], Qwen2-0.5B/7B-Instruct [39] and Qwen2.5-7B/14B-Instruct [40] as our language models. For brevity, we omit the “-Instruct” suffix in the remainder of this paper. We employed the most popular CLIP-ViT-L/14@336px[28] for the visual encoder. The connector is based on an MLP architecture with a hidden size of 5120. Regarding the language base vector, we carefully selected 19,200 high-frequency language tokens. Following the LLaVA [18] setup, we utilized LLaVA-CC3M-Pretrain-595K and LLaVA-v1.5-mix665k as the pretraining and instruction-tuning datasets, respectively, training for one epoch. Specific hyperparameter settings can be found in Appendix.
5.1.2 Benchmarks.
We evaluated our method on seven widely-used benchmarks: GQA [11], TextVQA [32], MME [43], MMBench [20], MMVeT [46], POPE [14], SciQA [30], covering a diverse range of vision-language understanding tasks. More details are in Appendix.
5.2 Experiment results
5.2.1 Same-architecture transfer
We first examine the effectiveness of LangBridge in the context of same-architecture transfer. This setting is of great significance for efficiently scaling up model size, as it allows leveraging knowledge learned on smaller, more readily trainable models to improve the performance of larger ones. As shown in Table 1, the LangBridge module, pre-trained or fine-tuned on a Qwen2-0.5B model, can be seamlessly transferred to Qwen2-7B and Qwen2.5-7B/14B models without requiring computationally expensive re-pretraining. The results demonstrate that this transfer not only maintains strong performance but also, in several instances, exceeds the performance of the baseline models trained directly on the target LLM size. For example, when transferring the Qwen2-0.5B-SFT-LangBridge connector to Qwen2.5-14B, we observe improvements in TextVQA (+1.14%) and MMVeT (+1.86%) compared to the baseline. Similarly, the Qwen2-0.5B-Pretrain-LangBridge connector improves SciQA compared to the Qwen2-7B baseline (+2.68%). These improvements, while modest, highlight LangBridge’s ability to effectively leverage and transfer learned representations, leading to enhanced performance even without further pre-training on the larger models.
5.2.2 Cross-architecture transfer
Table 4: Ablation study on hidden sizes. We compare different hidden sizes (5120 vs 1024) for LangBridge across various model configurations. Results show that our approach maintains strong performance even with reduced hidden sizes.
LLM Connector Hidden Size Benchmark Results
GQA TextVQA MME MMBench MMVeT POPE SciQA
LLaMA3-8B MLPs 4096 63.24 55.66 1736 71.0 31.0 87.1 78.61
- LangBridge 5120 64.00 (101.20%) 56.55 (101.60%) 1775 (102.25%) 72.6 (102.25%) 31.1 (100.32%) 87.7 (100.69%) 77.98 (99.20%)
- LangBridge 1024 63.73 (100.78%) 56.71 (101.89%) 1773 (102.13%) 70.5 (99.30%) 33.8 (109.03%) 87.4 (100.34%) 78.26 (99.55%)
Table 5: Ablation study on vocabulary sizes. We conduct experiments with different vocabulary sizes (19,200, 25,600, and 32,000) for LangBridge. Results demonstrate that a vocabulary size of 19,200 achieves optimal performance while maintaining minimal parameter count.
LLM Connector Vocab Size Benchmark Results
GQA TextVQA MME MMBench MMVeT POPE SciQA
Qwen2-7B Qwen2-0.5B-SFT- 19200 63.15 57.34 1904 71.0 31.6 88.3 79.25
- - Qwen2-0.5B-SFT-LangBridge 25600 63.13 (99.97%) 57.58 (100.42%) 1842 (96.74%) 71.8 (101.13%) 32.9 (104.11%) 87.9 (99.55%) 79.01 (99.70%)
- - Qwen2-0.5B-SFT-LangBridge 32000 63.11 (99.94%) 57.19 (99.74%) 1832 (96.22%) 72.7 (102.39%) 33.2 (105.06%) 88.6 (100.34%) 79.11 (99.82%)
Furthermore, we explore whether LangBridge demonstrates excellent performance in cross-architecture transfer scenarios (e.g., Qwen, LLaMA), specifically examining transfers from Qwen series models to LLaMA3 and vice versa. As shown in Table 2, LangBridge demonstrates a remarkable ability to maintain strong performance, retaining a high percentage of the baseline performance, exceeding 99% on many benchmarks. Moreover, the transfer from Qwen2-0.5B to Llama3-8B results in significant improvements, with performance gains observed across most benchmarks, especially MMVeT. In the reverse direction, transferring from LLaMA3-8B to Qwen models (7B to 14B), LangBridge maintains robust performance, consistently achieving above 99% of the baseline scores across most benchmarks. In several cases, it even surpasses the baseline performance, such as when transferring to Qwen2.5-7B, where the LLaMA3-8B-SFT-LangBridge connector improves MMVeT (+8.33%) and POPE (+1.92%). Similarly, when applied to Qwen2.5-14B, the connector achieves gains in GQA (+1.19%) and MMVeT (+1.59%). These results demonstrate LangBridge’s strong capability for cross-architecture transfer, effectively bridging different model families while maintaining or even enhancing performance across diverse multimodal tasks.
5.2.3 Results of normal setting
Having demonstrated LangBridge’s unique capabilities in same-architecture transfer and cross-architecture transfer, we now turn to its direct performance as a multimodal training paradigm compared to the standard MLP approach used in LLaVA. Table 3 demonstrates that LangBridge achieves performance comparable to the standard MLP adapter across a range of benchmarks. Notably, LangBridge consistently matches or exceeds the performance of MLPs on LLaMA3-8B for most metrics (GQA, TextVQA, MME, MMBench, MMVeT, and POPE, with SciQA slightly degrades). Furthermore, competitive results are also achieved on the smaller Qwen2.5-7B model, with improvements observed particularly in MMVeT.
These results indicate that LangBridge is not only an effective transfer learning technique but also a competitive alternative to standard MLP-based training paradigms for multimodal large language models.
5.2.4 Ablation Study
To thoroughly validate the effectiveness of our LangBridge across different architectural configurations, we first conduct ablation studies on the MLP hidden size, followed by an analysis of different vocab size.
The effect of hidden size. While our main experiments used a fixed hidden size of 5120 as described in Sec 5.1.1, we performed additional experiments with a reduced hidden size of 1024 to investigate the impact of model capacity. As shown in Table 4, when reducing the hidden size from 5120 to 1024, our LangBridge still maintains strong performance across different benchmarks. Specifically, LLaMA3-8B-LangBridge-1024 achieves gains on GQA (+0.78%), TextVQA (+1.89%), MME (+2.13%), MMVeT (+9.03%), and POPE (+0.34%) compared with the LLaMA3-8B-MLPs. This demonstrates the effectiveness of our LangBridge across different hidden sizes.
The effect of vocab size. Table 5 compares LangBridge’s performance with different vocabulary sizes. While the 25,600-vocab version achieves marginal improvements on 3 benchmarks (TextVQA, MMBench, MMVeT) and the 32,000-vocab version also shows slight gains on 3 metrics, both exhibit notable declines on multimodal benchmarks like MME (-6.22%/-7.56%). Given that the 19,200-vocab configuration maintains optimal overall performance with minimal parameters, we select it as our final design.
6 Conclusions
In this work, we explore how MLPs in LVLMs bridge the vision-language modality gap. Our investigation reveals that MLPs progressively learn to project visual embeddings into subspaces spanned by their corresponding text embeddings. Drawing inspiration from this insight, we propose LangBridge, a novel adapter that translates visual features by decomposing them into linear combinations of LLM’s vocabulary embeddings. This design effectively tackles a key challenge: the need for adapter retraining when switching between different LLM backbones. Extensive experiments show that LangBridge- an adapter pretrained on Qwen2-0.5B can be applied to larger models like LLaMA3-8B or Qwen2.5-14B maintaining competitive performance.
[…]
Supplementary Material
Overview
In this supplementary material, we present more dataset details and more experimental results that are not included in the main paper. The contents include:
  • A comprehensive introduction to evaluation benchmarks S-A.
  • Detailed training configurations and hyperparameters S-B.
  • Discussion of current limitations and future directions S-C.
S-A Evaluation Benchmarks
We evaluated our method on seven widely-used benchmarks, covering a diverse range of vision-language understanding tasks.
  • GQA [11]: Evaluates the model’s visual perception ability through open-ended questions.
  • TextVQA [32]: Tests the model’s ability to read and reason about text in images to answer questions, focusing on text-based visual reasoning.
  • ScienceQA [30]: Provides a set of multiple-choice science questions with images to test the model’s zero-shot generalization ability in scientific question answering.
  • MME [43]: A comprehensive evaluation of LVLMs across ten perception tasks (e.g., OCR and object recognition) and four cognitive tasks (e.g., commonsense reasoning, numerical computation, translation, and code reasoning).
  • MMBench [20]: A bilingual benchmark for evaluating LVLM’s multimodal understanding capabilities, consisting of approximately 3000 multiple-choice questions covering 20 ability dimensions. The Chinese version is called MMBench-CN.
  • MMVeT [46]: A challenging multimodal benchmark designed to evaluate vision-language models’ robustness and reliability. It focuses on testing fine-grained visual understanding, complex reasoning, and real-world application scenarios.
  • POPE [14]: Evaluates the model’s ability to identify specific objects in images, aiming to detect object-level hallucinations. It uses ”yes/no” questions based on object annotations, with 50% of queries targeting existing objects and 50% targeting non-existing objects, employing random, popular, and adversarial sampling strategies.
Through these comprehensive benchmarks, we systematically evaluated the model’s capabilities across diverse tasks, with particular emphasis on multimodal understanding, visual reasoning, hallucination detection, and real-world applicability. These benchmarks collectively provide a thorough assessment of the model’s strengths and potential areas for improvement.
S-B Training Details
As shown in Table S-1, we follow the same training recipes as LLaVA’s settings [19] for standard MLPs, except that we change the pretrain learning rate from 1e-3 to 2e-5 for LangBridge.
Table S-1: Training hyper-parameters
Hyper-parameter Value
batch size 256 (pretrain), 128 (finetune)
learning rate 1e-3 (pretrain), 2e-5 (finetune)
learning rate schedule cosine
learning rate warm-up ratio 0.03
weight decay 0
epoch 1
optimizer AdamW
float precision bfloat16
deepspeed configuration zero2 (pretrain), zero3 (finetune)
S-C Limitations
While our work demonstrates promising results in vision-language tasks, it has a notable limitation: we only focus on the visual modality. Modern multimodal systems often need to process various types of inputs beyond images, such as videos, audio, and 3D data. Future work could explore extending LangBridge to support these additional modalities, potentially enabling a more comprehensive multimodal understanding system.